Session: May 2001 (session 2)
Wavelength: 6.0 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Marco Morsiani at Sun Jun 3 23:44:38 METDST 2001) Success

Noto (P. Leto at Mon Jun 4 08:17:33 METDST 2001) Minor failures
Lost 16:03-16:07 and 18:30-19:00 for ACU fault. LSB power for BBC08 variable in the range 5000-65535

Torun (K.Borkowski at Mon Jun 4 09:14:28 METDST 2001) Minor failures
BBC5 and BBC8 out of order (very low levels) throughout the experiment. No other known problems.

Jodrell_Bank (Alastair Gunn at Mon Jun 4 18:32:41 METDST 2001) Success
No known problems.

Cambridge (Alastair Gunn at Mon Jun 4 18:33:06 METDST 2001) Success
Formatter re-set prior to this experiment due to earlier power failure. Offset may have changed. No other known problems.

Onsala (John Conway at Tue Jun 5 11:08:32 METDST 2001) Success
No known problems

Westerbork (Tony Foley at Tue Jun 5 16:02:54 METDST 2001) Minor failures
Phasecal stopped working. Otherwise believed OK

Effelsberg (A. Kraus at Wed Jun 6 08:35:56 METDST 2001) Success